Firm A and Firm B have the same assets but different capital structures. Suppose Firm A has no debt and Firm B has $1,000,000 debt with 10% interest. If the tax rate for the two firms is 30%, what is the amount of annual tax savings for Firm B compared to Firm A?
Answer #1

Annual Interest that needs to be paid by firm B = 1,000,000 * 0.1 = 100,000

Tax saving by firm B = 0.3 * 100,000 = 30,000

Therefore, $30,000 s the amount of annual tax savings for Firm B compared to Firm A.